OpenStudy (anonymous):

Which word in the sentence is the indirect object? Tonight, the Bensons will be showing us pictures of their trip to Berlin. A. us B. Berlin C. pictures D. trip

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

\[\ge \le \forall \subset =\exists \sqrt[_{?}]{?} answer A\]

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

How are you LaTexing?

OpenStudy (anonymous):

so classey

OpenStudy (anonymous):

wait do coller code work on heer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

[ff00ff]Nope

OpenStudy (anonymous):

No click equation

OpenStudy (anonymous):

nope thay dont lol

OpenStudy (anonymous):

\[Just click equation\]

OpenStudy (anonymous):

No, like \{\color{blue} like that.

OpenStudy (anonymous):

it dident work xD

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

nexttttttt